OpenStudy (anonymous):

find the perpendicular bisector of the segment joining (0,2) and (3,0)

OpenStudy (anonymous):

first find the slope of (0,2)(3,0) which is -2/3... then flip this, so you get 3/2 as the slope of the bisector.. now what?

OpenStudy (anonymous):

you have done the hard work. now find the midpoint of the line segment

OpenStudy (anonymous):

average the coordinates you get \[(\frac{3}{2},1)\]is the midpoint. use point-slope formula and you are done

OpenStudy (anonymous):

Now find the mod points of the segment \[(\frac{x_2 +x_1}{2}),(\frac{y_2 +y_1}{2})\]

OpenStudy (anonymous):

\[(\frac{x_2 +x_1}{2},\frac{y_2 +y_1}{2})\] edited

OpenStudy (anonymous):

yes, I know the mid points.. but.. okay so, 1= SLOPE OF WHAT (3/2) + b

OpenStudy (anonymous):

slope of the bisector which is.. 3/2.. which is the same as the x...?

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \[x_1=\frac{3}{2},y_1=1,m=\frac{3}{2}\]

OpenStudy (anonymous):

y=mx+b please :D

OpenStudy (anonymous):

\[y-1=\frac{3}{2}(x-\frac{3}{2})\] etc

OpenStudy (anonymous):

ok fine

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

\[y-1=\frac{3}{2}x-\frac{9}{4}\] \[y=\frac{3}{2}x-\frac{5}{4}\]

OpenStudy (anonymous):

and you think your brain hurts!

OpenStudy (anonymous):

yeah Private school precalc summer packs are no joke. Neither is my teacher. Prepare for a long year :)